OpenStudy (anonymous):

Medal to best help 1. If the legs of a right triangle have lengths of 6 inches and 8 inches, what is the length of the hypotenuse in inches? (1 point) 11 13 10*** 16 Is my answer correct? I starred the answer I think is correct.

OpenStudy (anonymous):

@amistre64 @hartnn @thomaster

undeadknight26 (undeadknight26):

i also think it is 10... a^2 + b^2 = c^2 (6*6) + (8*8) = 36 +64 = 100 sqrt100=10

undeadknight26 (undeadknight26):

hope i helped...

OpenStudy (anonymous):

Thank you!! Could you check a few other questions? I hope that doesn't bother you! @undeadknight26

undeadknight26 (undeadknight26):

uh i could try though...bu all i really did was google it and see what most people said honestly..

undeadknight26 (undeadknight26):

Is this one of your questions? If the hypotenuse is 41 inches, and one of the legs is 40 inches, what is the length of the other leg in inches? (1 point) 65 57 25 9

OpenStudy (anonymous):

Yes

undeadknight26 (undeadknight26):

Not sure if i am allowed to do this but this should help you out alot... http://answers.yahoo.com/question/index?qid=20121025091315AAmjwXQ

OpenStudy (anonymous):

Ok...well thank you :)

undeadknight26 (undeadknight26):

sorry i will get back to you in like 5 years when i learn this if those are wrong:D

OpenStudy (anonymous):

lol...ok